Please help solve this problem using the given equation and kindly write down all steps clearly.
media%2Ff33%2Ff33b9b47-ff94-4b13-9ff6-52
media%2Feae%2Feae93e3a-7e50-4bd2-ae1d-0c
Answer #1

Theory: We need to first determine the Cp of ideal meathane gas. Then we will us ethe integration equation to calculate the heat released.

Solution: Step1:

we know for an ideal gas, Cp-Cv =R where Cp is heat capacity at constant pressure.

Cv is heat capacity at constant volume and R is gas constant =8.314 J/molK

We also know that for ideal gas, Cv = degree of freedom/(2*R)

CH4 has 3 linear degree of freedom, 3 rotational degree of freedom and 9 vibrational degrees of freedom. So CH4 has total 15 degrees of freedom.

So Cv = 15/2*R

Cp = Cv +R = 15/2*R +R = 17/2*R = 8.5R

Step 2:

873.15cg -873.15 8.5H 533.15 H 533.15

Q = 8.5R{873.15-533.15] = 8.5*8.314*340 = 24027.46 J = 24.027 kJ

I hope i was very detailed in answering your query. Kindly let me know if you still have any further doubt. if you are satisfied then please rate me. Thanks for using Chegg.